OpenStudy (anonymous):

estimate the parameter, give the center, radius and the and indicate the positive orientation

OpenStudy (anonymous):

x=1-3sin4(pi)t, y=2+3cos4(pi)t, 0<=t=<1/2

OpenStudy (anonymous):

anybody?